OpenStudy (anonymous):

I NEED HELP!! I DON"T UNDERSTAND!!! Given that P = (7, 16) and Q = (-1, 7), find the component form and magnitude of qp

OpenStudy (anonymous):

@ganeshie8 @mathstudent55 @SolomonZelman

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@robtobey @tkhunny

OpenStudy (anonymous):

\(\vec{QP} = \vec{P} - \vec{Q} = (7+1, 16-7) = (8,9)\)

OpenStudy (anonymous):

Magnitude |\(\vec{QP}\)| = \(\sqrt{8^2 +9^2} = \sqrt {145}\)

OpenStudy (anonymous):

Does that help ?

OpenStudy (anonymous):

oh my gosh thank you sooooo much!!!

OpenStudy (anonymous):

You are welcome :)